# Hunting Primordial Black Hole Dark Matter in Lyman-$\alpha$ Forest

### v1

@@ IV Lyman-$\alpha$ Forest as a probe of IGM temperature
Lyman-$\alpha$ transitions occur in hydrogen atom when the electron is excited from the ground state ($n=1$) to the first excited state ($n=2$) and vice versa. Thus, whenever a photon with wavelength $\sim$1216 Å, which corresponds to the energy difference between the ground state and first excited state, encounter neutral hydrogen they are absorbed along the line of sight. Photons coming towards us from a distant quasar, encounter neutral hydrogen in IGM at different redshifts, and get absorbed. Due to the expansion of the Universe, these transitions lead to a series of absorption lines in the observed quasar spectrum. The resulting spectrum resembles a forest, and thus called Lyman-$\alpha$ forest. The Lyman-$\alpha$ absorption is quantified by the optical depth ($\tau$) along the line of sight, given by [136],
$\tau_{\mathrm{Ly}}(z)=1.3\Delta_{b}\left(\frac{x_{\mathrm{HI}}}{10^{-5}}\right)\left(\frac{1+z}{4}\right)^{3/2}\left(\frac{dv/dx}{H(z)/(1+z)}\right)^{-1},\phantom{xxx}$ (20)
where $z$ is the redshift of the gas at fixed density with a line of sight gradient, $dv/dx$ , $\Delta_{b}$ is the baryonic density in units of cosmic mean density, $\Delta_{b}=\frac{\rho_{b}}{\langle\rho_{b}\rangle}$, where $\rho_{b}$ is the baryonic density. In the above equation, $x_{\mathrm{HI}}$ is the neutral hydrogen fraction. Using the ‘fluctuating Gunn-Peterson approximation’ [137, 138], peculiar velocity gradient, $dv/dx\sim H(z)/(1+z)$.
@@
From measurements of both $T_{0}$ and $\gamma$ as a function of redshift, one can know about the temperature evolution of the IGM during and after hydrogen and helium reionization [88, 89, 90, 91, 92, 67]. High redshift quasar spectra are compared with mock Lyman-$\alpha$ forest generated by hydrodynamical simulations, to infer these parameters. Some of the widely used statistics are Voigt profile decomposition, flux power spectrum, probability distribution of flux [146, 142, 143, 147, 148, 149, 150, 71, 151, 152, 153, 154, 155, 66, 156, 65, 67, 157]. Most of these measurements use complementary methods. Below we briefly talk about the datasets that we use in this work:
@@
Lyman-$\alpha$ flux power spectrum probes a large scale starting from $\sim$ 500 kpc to $\sim$ 10 Mpc. The shape of the power spectrum and the small scale cutoffs help in breaking degeneracies between thermal parameters like $T_{0}$ and $\gamma$. Ref. [65] used high resolution quasar spectra from surveys like BOSS [158], XQ-100 [159], Keck/HIRES [160, 161], Magellan/MIKE [162] and compared the resulting power spectra with those obtained from hydrodynamical simulation [163], to obtain temperature measurements in the redshift range $5.21.8\,\textless\,z\,\textless\,$ < 5.2.
@@
Ref. [66] looked at high resolution spectra of 15 quasars obtained from Keck/HIRES [164] and VLT/UVES [165]. Comparing the flux power spectra from observational data with predictions from a suite of hydrodynamical simulations, they obtained the IGM temperature at mean density in the redshift range $5.24\,\textless\,z\,\textless\,$ < 5.2.
@@
At $z\,\textgreater\,5$, the Lyman-$\alpha$ optical depth increases with redshift. As a result, the transmitted Lyman-$\alpha$ flux becomes close to zero with few transmission spikes. These spikes correspond to underdense, highly ionized gas [166, 167, 168]. Ref. [67] investigated the dependence of these spike widths on the gas temperature. By comparing spectra from five QSOs (Quasi-Stellar Objects) at $z\,\textgreater\,6$ with hydrodynamical simulations, the authors obtained the gas temperature at mean density in the redshift range $5.85.4\,\textless\,z\,\textless\,$ < 5.8.
@@
In Fig. 2, we show our constraints on evaporating PBH DM. In the left panel we show our bounds on non-spinning PBH DM for ‘conservative’ (red line) and ‘photoheating’ (blue line) models. The deep shaded regions show the variation in the constraints when we vary the electron fraction limits for ‘Tanh’, ‘FlexKnot’, and ‘Trac’ models. The regions above the red and blue lines are ruled out by the datsets used in this work. For comparison we also show the combined previous limits on PBH DM from various different astrophysical and cosmological observables [43, 42, 45, 46, 47, 48, 49, 50, 51, 54, 52]. We also show our bounds when the spin parameter of PBH is, $a_{*}=0.9999$, in the right panel of Fig. 2. The corresponding previous limits on spinning PBH DM are also shown. As expected, we find that when we do not include any effects of reionization, we obtain less stringent constraints. But these constraints are completely robust to any astrophysical uncertainties related to the process of reionization. In the case of ‘photoheating’ model, both photoionization and photoheating are included, which results in stronger constraints on PBH DM. Since the ‘photoheating’ constraints are affected by photoheating during reionization, hence how reionization proceeds affects these constraints much more, and thus these constraints are represented by a wider band as compared to the ‘conservative’ constraints. In both non-spinning and spinning cases, we find that our limits are complementary to the limits present in the parameter space. Our limits are derived assuming a monochromatic mass distribution for PBH DM. We leave the exploration of the Lyman-$\alpha$ forest limits with other mass distributions (like log-normal distribution) for future work.
For some other constraints exploiting the evaporation of PBHs, the reader is urged to consult [169, 170, 46, 171, 172, 173, 174, 175, 176, 177, 49, 60, 59, 178, 179, 54, 62, 51, 52, 180]. QED corrections to Hawking radiation have been discussed in [181, 182, 183]. Constraints on evaporating PBH DM that take into account a possible quantum gravity effect, termed ‘memory burden effect’, can differ from those based on the standard Hawking radiation emission and are discussed in [184, 185, 186, 187, 188]. There have also been studies on PBH formation in alternate metrics [189, 190]. Heavier PBHs have also been constrained using various other techniques, a non-exhaustive assortment is [191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211]. We note in passing that PBHs in the mass window $10^{18}$ g to $10^{22}$ g have proven to be notoriously challenging to probe. Some papers that discuss various ideas for probing PBHs of these masses are [212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226].
@@ A.1 Method A
It may come off as baffling that how this spectrum could depend on $\tau$ and $m$, whose values, as remarked above, are arbitrary. The way to unpack this is to notice that the rate of total energy-injection per unit volume by decaying particle dark matter is given by
$\left(\frac{dE}{dVdt}\right)_{\mathrm{inj}}=\frac{\rho_{\chi}}{\tau},$ (26)
where $\rho_{\chi}$ is the density of dark matter and $\tau$ is the lifetime of the particle constituting dark matter. The R.H.S. of the above equation has to be multiplied by the spectrum given off by each particle’s decay. The $\tau$ factors clearly cancel. Since $\rho_{\chi}\propto m$, the $m$ factors also cancel off. Thus, it does not matter what values of $\rho$ and $m$ we use, as long as they are reasonable values for particle dark matter.
@@ A.2 Method B
The default energy binning in DarkHistory is different than BlackHawk. So the spectra from BlackHawk have to be properly binned to the default energy binning of DarkHistory. Next, in the pppc.py file, for a particular PBH mass, we input the ($d^{2}N/dEdt$) in photon and electron-positron channels. Note that the default Spectrum function uses ($dN/dE$) instead of ($d^{2}N/dEdt$). Thus, we simply use the same $\Delta$t timestep specified in the main.py file of the default DarkHistory code, to convert $d^{2}N/dEdt$ to $dN/dE$.

### v2

@@ IV Lyman-$\alpha$ Forest as a probe of IGM temperature
Lyman-$\alpha$ transitions occur in the hydrogen atom when the electron is excited from the ground state ($n=1$) to the first excited state ($n=2$) and vice versa. Thus, whenever photons with wavelength $\sim$1216 Å, which corresponds to the energy difference between the ground state and first excited state, encounter neutral hydrogen they are absorbed along the line of sight. Photons coming towards us from a distant quasar, encounter neutral hydrogen in IGM at different redshifts, and get absorbed. Due to the expansion of the Universe, these transitions lead to a series of absorption lines in the observed quasar spectrum. The resulting spectrum resembles a forest, and is thus called Lyman-$\alpha$ forest. The Lyman-$\alpha$ absorption is quantified by the optical depth ($\tau$) along the line of sight, given by [142],
$\tau_{\mathrm{Ly}}(z)=1.3\Delta_{b}\left(\frac{x_{\mathrm{HI}}}{10^{-5}}\right)\left(\frac{1+z}{4}\right)^{3/2}$
$\times\left(\frac{dv/dx}{H(z)/(1+z)}\right)^{-1},$ (20)
where $z$ is the redshift of the gas at fixed density with a line of sight gradient, $dv/dx$ , $\Delta_{b}$ is the baryonic density in units of cosmic mean density, $\Delta_{b}=\frac{\rho_{b}}{\langle\rho_{b}\rangle}$, $\rho_{b}$ being the baryonic density. In the above equation, $x_{\mathrm{HI}}$ is the neutral hydrogen fraction. Using the ‘fluctuating Gunn-Peterson approximation’ [143, 144], peculiar velocity gradient, $dv/dx\sim H(z)/(1+z)$.
@@
From measurements of both $T_{0}$ and $\gamma$ as a function of redshift, one can know about the temperature evolution of the IGM during and after hydrogen and helium reionization [94, 95, 96, 97, 98, 73]. High redshift quasar spectra are compared with mock Lyman-$\alpha$ forest bluespectra generated by hydrodynamical simulations, to infer these parameters. These inferences are based on the Voigt profile decomposition, flux power spectrum, probability distribution of flux. [152, 148, 149, 153, 154, 155, 156, 77, 157, 158, 159, 160, 161, 72, 162, 71, 73, 163]. Most of these measurements use complementary methods. Below we briefly talk about the datasets that we use in this work:
@@
Lyman-$\alpha$ flux power spectrum probes a large scale starting from $\sim$ 500 kpc to $\sim$ 10 Mpc. The shape of the power spectrum and the small scale cutoffs help in breaking degeneracies between thermal parameters like $T_{0}$ and $\gamma$. Ref. [71] used high resolution quasar spectra from surveys like BOSS [164], XQ-100 [165], Keck/HIRES [166, 167], Magellan/MIKE [168] and compared the resulting power spectra with those obtained from hydrodynamical simulation [169], to obtain temperature measurements in the redshift range $1.8\,\textless\,z\,\textless\,5.2$.
@@
Ref. [72] looked at high resolution spectra of 15 quasars obtained from Keck/HIRES [170] and VLT/UVES [171]. Comparing the flux power spectra from observational data with predictions from a suite of hydrodynamical simulations, they obtained the IGM temperature at mean density in the redshift range $4\,\textless\,z\,\textless\,5.2$.
@@
At $z\,\textgreater\,5$, the Lyman-$\alpha$ optical depth increases with redshift. As a result, the transmitted Lyman-$\alpha$ flux becomes close to zero with few transmission spikes. These spikes correspond to underdense, highly ionized gas [172, 173, 174]. Ref. [73] investigated the dependence of these spike widths on the gas temperature. By comparing spectra from five QSOs (Quasi-Stellar Objects) at $z\,\textgreater\,6$ with hydrodynamical simulations, the authors obtained the gas temperature at mean density in the redshift range $5.4\,\textless\,z\,\textless\,5.8$.
@@
In Fig. 2, we show our constraints on evaporating PBH DM. In the left panel we show our bounds on non-spinning PBH DM for ‘conservative’ (red line) and ‘photoheating’ (blue line) models. The deep shaded regions show the variation in the constraints when we vary the electron fraction limits for ‘Tanh’, ‘FlexKnot’, and ‘Trac’ models. The regions above the red and blue lines are ruled out by the datsets used in this work. For comparison we also show the combined previous limits on PBH DM from various different astrophysical and cosmological observables [49, 48, 51, 52, 53, 54, 55, 56, 57, 60, 58]. We also show our bounds when the spin parameter of PBH is, $a_{*}=0.9999$, in the right panel of Fig. 2. The corresponding previous limits on spinning PBH DM are also shown. As expected, we find that when we do not include any heating effects from reionization, we obtain less stringent constraints. But these constraints are completely robust to any astrophysical uncertainties related to the process of reionization. In the case of ‘photoheating’ model, both photoionization and photoheating are included, which results in stronger constraints on PBH DM. Since the ‘photoheating’ constraints are affected by photoheating during reionization, hence how reionization proceeds affects these constraints much more, and thus these constraints are represented by a wider band as compared to the ‘conservative’ constraints. In both non-spinning and spinning cases, we find that our limits are complementary to the limits present in the parameter space. Our limits are derived assuming a monochromatic mass distribution for PBH DM. We leave the exploration of the Lyman-$\alpha$ forest limits with other mass distributions (like log-normal distribution) for future work.
For some other constraints exploiting the evaporation of PBHs, the reader is urged to consult [175, 176, 52, 177, 178, 179, 180, 181, 182, 183, 55, 66, 65, 184, 185, 60, 68, 57, 58, 186]. QED corrections to Hawking radiation have been discussed in [187, 188, 189]. Constraints on evaporating PBH DM that take into account a possible quantum gravity effect, termed ‘memory burden effect’, can differ from those based on the standard Hawking radiation emission and are discussed in [190, 191, 192, 193, 194]. There have also been studies on PBH formation in alternate metrics [195, 196]. Heavier PBHs have also been constrained using various other techniques, a non-exhaustive assortment is [197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217]. We note in passing that PBHs in the mass window $10^{18}$ g to $10^{22}$ g have proven to be notoriously challenging to be probed. Some papers that discuss various ideas for probing PBHs of these masses are [218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233].
@@ A.1 Method A
It may come off as baffling that how this spectrum could depend on $\tau$ and $m$, whose values, as remarked above, are arbitrary. The way to unpack this is to notice that the rate of total number of dark matter particles decaying per unit volume by decaying particle dark matter is given by
$\left(\frac{dN_{\chi}}{dVdt}\right)_{\mathrm{inj}}=\frac{\rho_{\chi}}{m\tau},$ (26)
where $\rho_{\chi}$ is the density of dark matter, which is the same irrespective of the dark matter constituent, and $\tau$ is the lifetime of the particle constituting dark matter. The R.H.S. of the above equation has to be multiplied by the spectrum given off by each particle’s decay to get the total injection rate of photons per unit volume. The $\tau$ and $m$ factors clearly cancel off. Thus, it does not matter what values of $\tau$ and $m$ we use, as long as they are reasonable values for particle dark matter.
@@ A.2 Method B
[…]